40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 30 16:11:57 2016, ericq, Update, General, Vent Prep 
    Reply  Thu Jun 30 21:24:36 2016, ericq, Update, General, Vent Prep 
       Reply  Fri Jul 1 09:20:44 2016, Steve, Update, VAC, vent 78 vent78.pngvent78at250T.png
          Reply  Fri Jul 1 15:48:10 2016, Steve, Update, VAC, vent 78 at 710 Torr pd78at710T.png
             Reply  Wed Jul 6 10:58:07 2016, Steve, Update, VAC, vent 78 valve configuration CamberOpen.png
                Reply  Wed Jul 20 11:39:07 2016, Steve, Update, VAC, vent 78 day 20 vent78-d20.png
                   Reply  Mon Aug 8 10:35:12 2016, Steve, Update, PEM, vent 78 day 38 Vent38days.pngvent78d40RgaBg.png
                      Reply  Tue Oct 4 11:02:31 2016, Steve, Update, VAC, RGA is out of order lastBGscan.png
                         Reply  Fri Oct 7 10:29:57 2016, Steve, Update, VAC, RGA removed RGAremoved.png
                            Reply  Mon Oct 31 16:20:50 2016, Steve, Update, VAC, RGA is reinstalled 
                            Reply  Tue Nov 22 11:31:52 2016, Steve, Update, VAC, RGA is running again RGAisBack33d.pngpd80-560Hz-d33.pngafterRepair1d.png
                               Reply  Mon Nov 28 09:45:44 2016, Steve, Update, VAC, RGA scan at pumpdown day 40 RGAscan_d40.pngd2_&_d7scan.pngRGAd2.png
                                  Reply  Thu Dec 22 09:37:20 2016, Steve, Update, VAC, RGA scan at day 63 pd80-d63.pngpd80-580Hz-d63.png
                                     Reply  Tue Jan 10 14:24:09 2017, Steve, Update, VAC, RGA scan at day 82 pd80VNd82.png
                                        Reply  Mon Feb 27 15:33:10 2017, Steve, Update, VAC, RGA scan at day 130 rgascan@130d.png
          Reply  Wed Jul 6 08:34:00 2016, Steve, Update, VAC, RGA scan at day 639 RGAscan639d.pngpd78d640.png
          Reply  Mon Sep 12 17:54:24 2016, Steve, Update, VAC, pumpdown 79 
Message ID: 12233     Entry time: Thu Jun 30 16:11:57 2016     Reply to this: 12235
Author: ericq 
Type: Update 
Category: General 
Subject: Vent Prep 

I have updated the vent prep checklist on the wiki. Gautam and I did the following things from it:

  • Center all oplevs, transmon QPDs 
    • ETMX oplev has not been centered, since it's moving around so much, and we're going to immediately move the suspension anyways.
  • Align the arm cavities for IR and align the green lasers to the arms. 
    • AUX X Green was aligned while the X arm was well aligned. Soon thereafter, ETMX wandered away, but the green will remain a good reference
  • Update the SUS Dritmon values 
  • Reconcile all SDF differences 
  • Reduce input power to no more than 100mW by adjusting wave plate+PBS setup on the PSL table BEFORE the PMC. (Using the WP + PBS that already exist after the laser.) 

  • Replace 10% BS before MC REFL PD with Y1 mirror and lock MC at low power.
    • I don't think we've vented since the most recent slew of changes to the IMC servo, so its not surprising that the current low power scripts don't work. I'm working on locking the IMC, but this does not prevent us from initiating the vent tomorrow.

 

The following bullets have not yet been executed:

  • Close shutter of PSL-IR and green shutters at the ends 
  • Make sure the jam nuts are protecting bellows 
  • Check crane functionality & cleanliness 

  • Turn off HV into vacuum: OMC is not wired this time 
  • Particle count must be under 10,000 counts / cf min for 0.5 micron 
  • Check all metal window covers are on. 
  • Check 5 cylinders (24 cft size) of instrument grade air, called Alfa Gas 1 in stock.
ELOG V3.1.3-